OpenStudy (anonymous):

how do i factor the following 9-x2

OpenStudy (anonymous):

how do i factor 9-x2?

OpenStudy (aroub):

(x-3)(x+3)

OpenStudy (anonymous):

use the formula \[a ^{2}-b ^{2}=(a+b)(a-b)\]

OpenStudy (aroub):

do u want the rule (formula)??

OpenStudy (aroub):

oh lol look at mashys

OpenStudy (anonymous):

never give direct answers, that ll spoil the kids :P

OpenStudy (aroub):

yes i know .... lol! :D

OpenStudy (anonymous):

i just don't get any of this stuff, but algebra and trig is required for me to get my degree

OpenStudy (anonymous):

4x2+81-36x

OpenStudy (anonymous):

??

OpenStudy (anonymous):

9-x2 = (3-x)(3+x)